ESG Recruitment Proposal March 13, 2001

ESG Recruitment Proposal

Committee in Charge of Doing the Obvious (C.I.C.D.O.)

March 13, 2001

1 The Problem

This last term, ESG tried to achieve the ideal of accepting into the program anyone interested in joining. The results of this were disappointing at best. The increased enrollment of ESG noticeably increased the difficulty in scheduling classes, increased class sizes, and increased the demands on and for ESG instructors somewhat beyond their capacity. Interestingly, the amount of general community activity and the number of people dedicated to the program did not seem to increase accordingly.

ESG claims to be significantly different from the mainstream of MIT in its approach to and attitude toward academics. Our approach involves the more active role of students in the learning process, taking charge of the direction of their education and progressing through material at their own rate. Because learning need no longer based on the traditional class structure under this approach, ESG has developed an attitude towards learning based on the belief that undergraduate education should not be limited to subjects which prepare a student for the work force or advance his or her career. Small class sizes are just a means associated with these differences, because they allow students to have more effect on the direction of their classes and make classes more closely resemble seminars rather than the fire hose sources of required curriculum.

Many freshmen enter ESG unaware of these differences or expecting to take advantage of the small class sizes without accepting change in approach and attitude that ESG requires of them. There are many aspects which make ESG unique and different from the mainstream, and many of the people who join the program are not interested in these aspects. If these people can be discouraged from joining we may be able to achieve the original ideal of accepting all applicants without the effects observed this term and with enrollment better matched to our capacity.

3 DISCUSSION ESG Recruitment Proposal

March 13, 2001

2 The Solution

To do this, ESG must make its different flavor of academics clear and somewhat formal. For ESG to succeed best, no students may be exempt from participating in this style of learning. A requirement to this effect, which appeals to people seeking the differences in learning while discouraging those who are not will make considerable progress in helping students self-select themselves for ESG. This requirement should be for something that students interested in ESG’s approach to academics are already inclined to do, while reflecting ESG’s attitude for the good in diversity of education. In essence, it should be a requirement which places no real additional constrains on many students.

Currently students in ESG are required only enroll for two classes, and encouraged to enroll in three. In addition, under this proposal, they must choose to do at least one of the following:

A) Enroll in a new, self-structured research seminar. Students interested in this option may take the seminar for 6 units of credit or without credit (if they already have a 6 unit seminar they want to keep). The are divided into groups of about 5 under a single advisor/instructor. At the beginning of the semester, each student selects a subject he or she is interest in but does not know much about. Over the course of the semester as students learn about their topics through independent work, the instructor will keep track of their progress. The only requirement is that at the end of the semester the students use what they have learned to give their own single meeting seminar on the topic.

B) Participate in some ESG study group or not-for-credit seminar. The seminar need not be academic, but it must meet at least once per week. At the beginning of the semester ESG upperclassmen and staff should organizes at least 5 study groups for students interesting in this option. Examples for this are the Godel, Escher, Bach study group, the Experimental Singing Group, and the Mathematical Topics seminar.

C) Work on a large project or organize something to improve the ESG community. Examples include painting murals, organizing trips, cooking Friday lunch, sorting books, and fixing up the ESG area.

3 Discussion

The first possibility of a research seminar provides students who are interested in pursuing entirely self-directed learning an environment to do so for credit. The final seminar will improve the ESG community with a series of interesting topics to discuss late in the semester.

The second option will appeal to students who are interested in learning outside of their regular curriculum, and will increase the attendance of the various study groups.

The third option is very similar to the current “Community Activities” requirement, but 4 NEW RECRUITMENT STRATEGIES ESG Recruitment Proposal

March 13, 2001

places the emphasis on persistent projects which will enrich the ESG community for future years.

The above options are meant to be broad enough to allow anyone interested in ESG a way of satisfying the requirement without doing anything they would not be inclined to doing anyway.

4 New Recruitment Strategies

One effect of these new requirements for ESG will be decreased enrollment. Although the people lost will hopefully be ones who would get no more out of ESG than they would out of the mainstream, this is an opportunity to admit more of the kind of people who would get a lot out of our approach. Therefore, another important change for ESG must come in its recruitment strategies.

CICDO has come up with two additional rush items to attract freshmen:

A) Give freshmen information about ESG at the “Core Blitz: Academic Options” meeting during orientation.

B) Put up a bulletin board about ESG in a prominent location, such as on the infinite corridor or the student center, or a drop poster in Lobby 7, during Rush.